# MDPs with a State Sensing Cost

### v1

@@ 4 Solving $\mathcal{M}_{k}$ with Provable Guarantees
In this section, we describe conditions and approaches that enable provable optimality/suboptimality guarantees for the MDP $\mathcal{M}_{k}.$ First, we provide a sufficient condition for always sensing to be optimal. Subsequently, we analyse a truncated version of $\mathcal{M}_{k}$ and provide sufficient conditions for the optimal solution of this (finite) truncated MDP to agree with that of $\mathcal{M}_{k}.$
@@ 4.1 Optimality of always sensing
The threshold on state sensing cost in Theorem 4.1 may be interpreted as follows. It is strictly positive if and only if, for any action $a_{1}$ taken in any root state $j,$ there does not exist an action $a_{2}$ that is optimal in the baseline MDP on all states that lie in the belief support.
@@ 4.2 Analysis via state space truncation
We now consider a class of finite MDPs obtained via state space truncation of $\mathcal{M}_{k}.$ Specifically, the truncation is parameterized by $n\geq 0,$ the maximum number of consecutive blind actions the agent is permitted to take.
Formally, the truncated MDP, denoted by $\mathcal{M}_{k,n}$ is defined as follows. The state space is given by
@@
Note that for $n\geq 1,$ $\mathcal{S}=\mathcal{S}_{0}\subset\mathcal{S}_{n}\subset\mathcal{S}_{n+1}\subset\mathcal{S}_{\infty}.$ We find it convenient to categorize the states of $\mathcal{S}_{n}$ into ‘layers’ as follows: Let $\mathcal{L}_{0}$ be the set of root states (the $0^{\text{th}}$ layer), which correspond to the states where the agent knows its current state precisely. Next, we define $\mathcal{L}^{j}_{m}$ as the set of ‘descendants’ of the root state $j$ in the $m^{\text{th}}$ layer, i.e., the set of states corresponding to playing $m$ successive blind steps starting from the root state $j$. More formally,
$\mathcal{L}^{j}_{m}=\{\tilde{s}\in\mathcal{S}_{\infty}\mid\tilde{s}=(j,a_{1},\ldots,a_{m}),\text{ where }a_{1},\ldots,a_{m}\in A\}.$
@@
The action space $A_{n}$ of $\mathcal{M}_{k,n}$ is simply $A_{\infty}$, whereas the transition function $\mathcal{T}_{n}:\mathcal{S}_{n}\times A_{n}\times\mathcal{S}_{n}\rightarrow\mathbb{R}$ and the cost function $\mathcal{C}_{n}:\mathcal{S}_{n}\times A_{n}\rightarrow\mathbb{R}$ are given by:
$\mathcal{T}_{n}(\tilde{s}_{1},(a,\cdot),\tilde{s}_{2})$ =$=\begin{cases}\mathcal{T}_{\infty}(\tilde{s}_{1},(a,\cdot),\tilde{s}_{2}),&\text{if }\tilde{s}_{1}\notin\mathcal{L}_{n}\\ \mathcal{T}_{\infty}(\tilde{s}_{1},(a,sense),\tilde{s}_{2}),&\text{otherwise}\end{cases}$
$\mathcal{C}_{n}(\tilde{s},(a,\cdot))$ =$=\begin{cases}\mathcal{C}_{\infty}(\tilde{s},(a,\cdot)),&\text{if}\tilde{s}\notin\mathcal{L}_{n}\\ \mathcal{C}_{\infty}(\tilde{s},(a,sense)),&\text{otherwise}\end{cases}$
@@
Since $\mathcal{M}_{k,n}$ is a finite MDP, it admits an exact computation of its optimal policy $\pi^{*}_{\mathcal{M}_{k,n}}$ and optimal value function $V^{*}_{\mathcal{M}_{k,n}}.$ Of course, the complexity of this computation grows exponentially in $n,$ so this is only feasible for small values of $n.$
In the remainder of this section, we relate the solutions of the (finite, and therefore ‘tractable’) truncated MDPs $\{\mathcal{M}_{k,n}\}$ to one another, and to the solution of $\mathcal{M}_{k}.$
@@
Note that using the above result, one can determine a suitable truncation depth $N$ given a suboptimality tolerance, without having to solve $\mathcal{M}_{k,N}$ first. (A more refined suboptimality bound, expressed in terms of the solution of $\mathcal{M}_{k,N}$ is provided later in Theorem 4.4.)
Our next result provides a necessary and sufficient condition for the optimal policy for $\mathcal{M}_{k,n}$ to also be optimal for $\mathcal{M}_{k,n+1}.$ For $s\in\mathcal{S}$ and $a_{1},a_{2},\ldots,a_{m}\in A$, define
@@
Fix $N\geq 0.$
$V^{*}_{\mathcal{M}_{k,N}}(j)=V^{*}_{\mathcal{M}_{k,N+1}}(j)$ for all root states $j\in\mathcal{S}$ if and only if
@@
Note that $V^{*}_{\mathcal{M}_{k,N}}(j)=V^{*}_{\mathcal{M}_{k,N+1}}(j)$ for all $j\in\mathcal{S}$ implies that the optimal stationary policy $\pi^{*}_{M_{k,N}}$ for $\mathcal{M}_{k,N}$ is also optimal for $\mathcal{M}_{k,N+1}$ when starting in a root state (i.e., with knowledge of the starting state). However, it is important to note that this condition does not imply that $\pi^{*}_{M_{k,N}}$ is optimal for $\mathcal{M}_{k}$ when starting in a root state. Indeed, we provide a counter-example in Section 5 to demonstrate this. One needs a stricter condition on $\mathcal{M}_{k,N}$ to conclude that $\pi^{*}_{M_{k,N}}$ is optimal for $\mathcal{M}_{k};$ this is the focus of our next result.
@@
then the optimal stationary policy $\pi^{*}_{\mathcal{M}_{k,N}}$ of $\mathcal{M}_{k,N}$ is optimal for $\mathcal{M}_{k}$ when starting at any root state.
If (1) does not hold,
@@
Theorem 4.4 provides a sufficient condition (1) for the optimal policy for $\mathcal{M}_{k,n}$ to also be optimal for $\mathcal{M}_{k}$ (assuming the starting state is a root state). Even if this condition is violated, Theorem 4.4 provides a computable upper bound on the suboptimality of the policy $\pi^{*}_{\mathcal{M}_{k,N}}$ on $\mathcal{M}_{k}.$ Thus, Theorem 4.4 suggests a recipe for computing an optimal/near-optimal policy for $\mathcal{M}_{k}$: Iteratively solve $\mathcal{M}_{k,N}$ for increasing $N,$ until either (i) the condition (4.4) is satisfied, in which case the optimal policy just computed is also optimal for $\mathcal{M}_{k},$ or (ii) the suboptimality bound $\epsilon_{N}$ is acceptably small. [^2] In fact, the following stronger statement follows from the proof of Theorem 4.4: If, for any root state $j,$ it holds that $Z(i)+\alpha^{N+1}V_{AS;0}(i)\geq V^{*}_{\mathcal{M}_{k,N}}(j)\quad\forall\ i\in\mathcal{L}^{j}_{N+1},$ (3) then the optimal policy for $\mathcal{M}_{k}$ takes at most $N$ consecutive blind steps starting from $j.$ Thus, if (3) is satisfied at certain root states, one does not need to explore depths $N+1$ and beyond at these root states.
Lemma 4.5.
@@ 5 Numerical Case Studies
In this section, we present numerical experiments that validate and also complement the analytical results in the preceding sections. We also benchmark the proposed approaches against the state of the art.
Counter-example related to Lemma 4.3: We begin with an example that demonstrates that starting at a root state, if a certain policy $\pi$ is optimal for $\mathcal{M}_{k,N}$ as well as $\mathcal{M}_{k,N+1},$ that does not guaranteee that the $\pi$ is also optimal for $\mathcal{M}_{k}.$ Consider the two-state two-action baseline MDP shown in Figure 2, with sensing cost $k=0.005$ and discount factor $\alpha=1/2.$
010.720.720.720.72, 0.0660.0660.0660.0660.9340.9340.9340.934, 0.5020.5020.5020.5020.280.280.280.28, $0.0660.066$, 0.5020.5020.5020.5020.690.690.690.69, 0.290.290.290.290.4810.4810.4810.481, 0.410.410.410.410.310.310.310.31, 0.290.290.290.290.5190.5190.5190.519, $0.41$
Interestingly, we find in this example that the criterion for Theorem 4.4 (see (1)) is satisfied for root state 00 at $N=2$. Therefore, it is clear at that point that the optimal policy for root state 00 will take a maximal of $2$ blind steps, and we can restrict our search for the optimal policy starting at 00 until the $2^{\text{nd}}$ layer. The same criterion is not satisfied for root state $1$ for $N\leq 6.$
$N$ $\pi^{*}_{\mathcal{M}_{k,N}}(0)$ $V^{*}_{\mathcal{M}_{k,N}}(0)$ $\pi^{*}_{\mathcal{M}_{k,N}}(1)$ $V^{*}_{\mathcal{M}_{k,N}}(1)$
0 R $0.367061$ B $0.6796465$
1 R $0.367061$ B $0.6796465$
2 R $0.367061$ B $0.6796465$
3 R $0.367061$ B $0.6796465$
4 R $0.36703456$ BRRRR $0.67958256$
5 R $0.367029$ BRRRRR $0.6795691$
6 R $0.3670226$ BRRRRRR $0.6795541$
@@
Another 2-State 2-Action example: We now consider another 2-state 2-action baseline MDP example, as shown in Figure 3. We evaluate the above MDP on sensing costs $(a)\ k=0.01$ and $(b)\ k=0.25$.
@@
$\bullet$ Applying Theorem 4.1 on the above MDP gives a sensing cost threshold of $0.05$. Hence for case $(a),$ always sensing is optimal and hence the optimal policy for state 00 is $(R,sense)$ and for $1$ is $(B,sense)$.
@@
$\bullet$ For case $(b),$ we find that the optimal policy for both the root states remains unchanged after $N\geq 2$. It turns out that the conditions of Lemma 4.3 hold for $N=2$. However, the condition of Theorem 4.4 is satisfied for $N=4$ which suggests that the optimal policy should remain unchanged for $N\geq 4;$ see Figure 4.
$\bullet$ For case $(b),$ Optimal policies for $\mathcal{M}_{k,N}$ outperform our heuristic algorithm (see Section 3) for $N\geq 1;$ see Figure 4.
Note: For computing the lower bound on $V^{*}_{\mathcal{M}_{m}}$ in Figure 4, and in the remaining figures of this section, we have used the following bound: For any root state $j,$
$V^{*}_{\mathcal{M}_{k}}(j)\geq\min\Bigl{\{}\min_{i\in\mathcal{L}^{j}_{N+1}}\left(Z(i)+\alpha^{N+1}V_{AS;0}(i)\right),V^{*}_{\mathcal{M}_{k,N}}(j)\Bigr{.}\\ \Bigl{.}-\alpha\max_{s\in\mathcal{S}\setminus\{j\}}\left[V^{*}_{\mathcal{M}_{k,N}}(s)-\min_{r\in\mathcal{L}^{s}_{N+1}}\left(Z(r)+\alpha^{N+1}V_{AS;0}(r)\right)\right]^{+}\Bigr{\}}.$
Here, $[x]^{+}$ denotes the positive part of $x.$
010.70.70.70.7, 11110.30.30.30.3, 000.90.90.90.9, 000.20.20.20.2, 11110.70.70.70.7, 000.30.30.30.3, 11110.80.80.80.8, 11110.10.10.10.1, 00
@@
Figure 3: A two-state two-action baseline MDP with actions $\{Red,Blue\}$ and and $\alpha=0.5$
Figure 4: Applying Theorem 4.4 on MDP in Fig 3; $k=0.25$
Benchmarking Results
@@
Finally, we evaluated the performance of the proposed SPI heuristic against the ATM heuristic proposed in (Krale et al., 2023) (both self-implemented in Python), alongside several widely-used general-purpose offline POMDP planning algorithms, including SARSOP (Kurniawati et al., 2008), Fast Informed Bound (FIB) (Hauskrecht, 2000; Kochenderfer et al., 2022), using the POMDPs.jl (Egorov et al., 2017) ecosystem of Julia packages across various tasks in the Gymnasium environment (Towers et al., 2024). Algorithms such as Incremental Pruning (Cassandra et al., 1997) and PBVI (Pineau et al., 2003) exhibited prohibitive runtimes, failing to produce effective policies even for small state spaces, while other methods like Q-MDP (Littman et al., 1995) were consistently outperformed by our approach. Even popular online planning algorithms like PO-UCT (Silver & Veness, 2010) required significantly higher planning times to generate competitive policies
Experimental Setup: Experiments were conducted on a MacBook Air with an Apple M3 chip and 16GB of memory. Hyperparameters were set according to the defaults in POMDPs.jl, with adjustments clearly stated and made to ensure comparable performance or runtime. POMDPs were initialized with a uniform distribution over the starting states. For SARSOP, reward and sensing values were scaled by a factor of 1000 to match performance on the Frozen-Lake task, and policy computation time was increased from 1s to 100s for the Taxi task.
Frozen Lake ($|\mathcal{S}|=16/64\ \&\ |A|=4$) : In this task, the agent navigates across a frozen lake and receives a reward of $+1$ upon reaching the goal state, with a discount factor of $0.9$ (Towers et al., 2024). We consider the slippery case, where the agent moves in the intended direction with a probability of $\frac{1}{3}$, and in one of the two perpendicular directions with equal probability of $\frac{1}{3}$ for each. The experiments use the default 4x4 and 8x8 grid configurations and a customized, challenging 4x4 version; see Table 2. Our results are summarized in Table 3. Note: We state rewards rather that costs, averaged uniformly over all initial states; the last column shows the maximum computation time recorded across the different sensing cost choices.
@@
Stochastic Taxi ($|\mathcal{S}|=500\ \&\ |A|=6$) : In this environment, the agent must pick up passengers and drop them off at the desired locations in a 5x5 grid world. We use the noisy version described in (Dietterich, 1999), where each of the four navigation actions moves the agent in the intended direction with a probability of 0.8, and in one of the two perpendicular directions with an equal probability of 0.1 for each. The agent receives reward +20 for delivering a passenger, -10 for illegal “pickup” and “drop-off” actions, and -1 per step unless another reward is triggered, with a discount factor of 0.95. The initial state is uniformly sampled from 300 valid states where the passenger is neither at their destination nor inside the taxi. Our results for this example are summarized in Table 4.
S F F F
F H F H
F F F H
H F F G
F H S F
F G H F
F H H F
F F F F
S F F F F F F F
F F F F F F F F
F F F H F F F F
F F F F F H F F
F F F H F F F F
F H H F F F H F
F H F F H F H F
F F F H F F F G
Table 2: Default and custom-hard 4x4 grids with default 8x8 grid.
We see that in nearly all cases, the proposed SPI heuristic outperforms the ATM heuristic as well as FIB; we were unable to run FIB on the Stochastic Taxi model with 500 (root) states. SPI and SARSOP are comparable in performance, with either approach outperforming the other in some examples. However, note that SPI is significantly better than SARSOP on the Stochastic Taxi model. Finally, we note that the performance of SARSOP is highly sensitive to the tuning of several ‘hard to interpret’ hyperparameters; in contrast, SPI has only two (easily interpretable) hyperparameters.
Scenario Value Function (Rewards) for Sensing Costs Time (s)
0.001 0.005 0.01 0.05
Frozen-Lake 4x4 (Default)
   SPI $\mathbf{62.42}$ $\mathbf{36.53}$ $20.99$ $\mathbf{23.08}$ $0.4$
   $\pi_{\text{ATM}}$ $\mathbf{62.42}$ $36.52$ $6.72$ $16.57$ $0.04$
   $V^{*}_{\mathcal{M}_{k,3}}$ $\mathbf{62.42}$ $\mathbf{36.53}$ $20.47$ $-28.75$ $11.5$
   SARSOP $\mathbf{62.42}$ $\mathbf{36.53}$ $20.79$ $\mathbf{23.08}$ $1.7$
   FIB ($1000$ iter) $\mathbf{62.42}$ $31.43$ $\mathbf{23.08}$ $\mathbf{23.08}$ $8.3$
Frozen-Lake 4x4 (Hard)
   SPI $\mathbf{8.95}$ $\mathbf{3.69}$ $\mathbf{1.47}$ $1.35$ $0.4$
   $\pi_{\text{ATM}}$ $8.41$ 00 00 00 $0.04$
@@ F F F H F F F G
   SARSOP $\mathbf{8.95}$ $3.66$ $1.34$ $\mathbf{1.44}$ $1.6$
   FIB ($1000$ iter) $4.44$ 00 00 00 $9.5$
Frozen-Lake 8x8
   SPI $\mathbf{3.53}$ $3.33$ $3.33$ $3.33$ $3$
@@
Table 3: Performance comparison across different scenarios with varying sensing costs; rewards are multiplicatively scaled by $10^{3}$.
Taxi Sensing Costs
0.1 0.5 1 5
SPI                 Value $\mathbf{0.911}$ $\mathbf{-0.306}$ $\mathbf{-1.598}$ $\mathbf{-9.778}$
                      Time (s) 25.2 19.7 17.8 117.8
${\pi_{\text{ATM}}}$               Value $0.908$ $-0.359$ $-2.761$ $-19.664$
                      Time (s) 2 1.7 1.8 38.4
SARSOP ($100$s) Value $-22$ $-1.2$ $-2.705$ $-20$
Table 4: Performance Comparison of Average Cumulative Reward for Stochastic Taxi
6 Concluding remarks

### v3

@@ 4 Solving $\mathcal{M}_{k}$ with Provable Guarantees
In this section, we establish conditions and methods that yield provable optimality and suboptimality guarantees for any policy $\pi$ of MDP $\mathcal{M}_{k}$. We first present a sufficient condition under which always sensing is optimal. We then analyze a truncated version of $\mathcal{M}_{k}$ to derive a lower bound on its optimal value function, which in turn provides a computable bound on the suboptimality gap of any policy.
@@ 4.1 Optimality of always sensing
The threshold on state sensing cost in Theorem 4.1 is strictly positive if and only if, for any action $a_{1}$ taken in any root state $j,$ there does not exist an action $a_{2}$ that is optimal in the baseline MDP on all states that lie in the belief support.
@@ 4.2 Analysis via state space truncation
We now consider a class of finite MDPs obtained via state space truncation of $\mathcal{M}_{k}.$ Specifically, the truncation is parameterized by $n\geq 0,$ the maximum number of consecutive blind actions the agent is permitted to take. Formally, the truncated MDP, denoted by $\mathcal{M}_{k,n}$ is defined as follows. The state space is given by
@@
For $n\geq 1$, this yields $\mathcal{S}=\mathcal{S}_{0}\subset\mathcal{S}_{n}\subset\mathcal{S}_{n+1}\subset\mathcal{S}_{\infty}.$ It is convenient to categorize the states of $\mathcal{S}_{n}$ into ‘layers’ as follows: Let $\mathcal{L}_{0}$ be the set of root states (the $0^{\text{th}}$ layer), where the agent knows its current state precisely. Next, we define $\mathcal{L}^{j}_{m}$ as the set of ‘descendants’ of the root state $j$ in the $m^{\text{th}}$ layer, i.e., the set of states corresponding to playing $m$ successive blind steps starting from the root state $j$. More formally,
$\mathcal{L}^{j}_{m}=\hskip-1.0pt\{\tilde{s}\in\mathcal{S}_{\infty}\hskip-2.0pt\mid\hskip-1.0pt\tilde{s}=(j,a_{1},\ldots,a_{m}),\hskip-3.0pt\text{ where }a_{1},\ldots,a_{m}\in A\}.$
@@
The action space $\mathcal{A}_{n}$ of $\mathcal{M}_{k,n}$ is simply $\mathcal{A}_{\infty}$, whereas the transition function $\mathcal{T}_{n}:\mathcal{S}_{n}\times\mathcal{A}_{n}\times\mathcal{S}_{n}\rightarrow\mathbb{R}$ and the cost function $\mathcal{C}_{n}:\mathcal{S}_{n}\times\mathcal{A}_{n}\rightarrow\mathbb{R}$ are given by:
$\mathcal{T}_{n}(\tilde{s}_{1},(a,\cdot),\tilde{s}_{2})$ $=\begin{cases}\mathcal{T}_{\infty}(\tilde{s}_{1},(a,\cdot),\tilde{s}_{2}),&\text{if }\tilde{s}_{1}\notin\mathcal{L}_{n}\\ \mathcal{T}_{\infty}(\tilde{s}_{1},(a,sense),\tilde{s}_{2}),&\text{otherwise}\end{cases}$
$\mathcal{C}_{n}(\tilde{s},(a,\cdot))$ $=\begin{cases}\mathcal{C}_{\infty}(\tilde{s},(a,\cdot)),&\text{if }\tilde{s}\notin\mathcal{L}_{n}\\ \mathcal{C}_{\infty}(\tilde{s},(a,sense)),&\text{otherwise}\end{cases}$
@@
Since $\mathcal{M}_{k,n}$ is a finite MDP, it admits an exact computation of its optimal policy $\pi^{*}_{\mathcal{M}_{k,n}}$ and optimal value function $V^{*}_{\mathcal{M}_{k,n}}.$ Of course, the complexity of this computation grows exponentially in $n,$ so this is only feasible for small values of $n.$ In the remainder of this section, we relate the solutions of the (finite, and therefore ‘tractable’) truncated MDPs $\{\mathcal{M}_{k,n}\}$ to one another, and to the solution of $\mathcal{M}_{k}.$
@@
Note that using the above result, one can determine a suitable truncation depth $N$ given a suboptimality tolerance, without having to solve $\mathcal{M}_{k,N}$ first. Moreover, solving $\mathcal{M}_{k,N}$ yields a computable lower bound on the optimal value function $V^{*}_{\mathcal{M}_{k}}$ of $\mathcal{M}_{k}$, which can then be used to bound the suboptimality gap of any policy. (A sharper bound, expressed in terms of the solution of $\mathcal{M}_{k,N}$ is provided later in Theorem 4.4.)
Our next result gives a necessary and sufficient condition for the optimal policy for $\mathcal{M}_{k,n}$ to also be optimal for $\mathcal{M}_{k,n+1}.$ For $s\in\mathcal{S}$ and $a_{1},a_{2},\ldots,a_{m}\in A$, define
@@
Fix $N\geq 0.$ $V^{*}_{\mathcal{M}_{k,N}}(j)=V^{*}_{\mathcal{M}_{k,N+1}}(j)$ for all root states $j\in\mathcal{S}$ if and only if
@@
Note that $V^{*}_{\mathcal{M}_{k,N}}(j)=V^{*}_{\mathcal{M}_{k,N+1}}(j)$ for all $j\in\mathcal{S}$ implies that the optimal stationary policy $\pi^{*}_{M_{k,N}}$ for $\mathcal{M}_{k,N}$ is also optimal for $\mathcal{M}_{k,N+1}$ when starting in a root state (i.e., with knowledge of the starting state). However, this condition does not imply that $\pi^{*}_{\mathcal{M}_{k,N}}$ is optimal for $\mathcal{M}_{k}$ when starting in a root state, as shown by a counterexample in Appendix E.2. One needs a stricter condition on $\mathcal{M}_{k,N}$ to conclude that $\pi^{*}_{\mathcal{M}_{k,N}}$ is optimal for $\mathcal{M}_{k};$ this is the focus of our next result.
@@
then the optimal stationary policy $\pi^{*}_{\mathcal{M}_{k,N}}$ of $\mathcal{M}_{k,N}$ is optimal for $\mathcal{M}_{k}$ when starting at any root state. If (1) does not hold,
@@
Theorem 4.4 provides a sufficient condition (1) for the optimal policy for $\mathcal{M}_{k,n}$ to also be optimal for $\mathcal{M}_{k}$ (assuming the starting state is a root state). Even if this condition is violated, Theorem 4.4 provides a computable upper bound on the suboptimality of the policy $\pi^{*}_{\mathcal{M}_{k,N}}$ on $\mathcal{M}_{k}.$ Thus, Theorem 4.4 suggests a recipe for computing an optimal/near-optimal policy for $\mathcal{M}_{k}$: Iteratively solve $\mathcal{M}_{k,N}$ for increasing $N,$ until either (i) the condition (4.4) is satisfied, in which case the optimal policy just computed is also optimal for $\mathcal{M}_{k},$ or (ii) the suboptimality bound $\epsilon_{N}$ is acceptably small. Moreover, analogous to Theorem 4.2, one can also derive a (stronger) lower bound on the optimal value function $V^{*}_{\mathcal{M}_{k}}$ of $\mathcal{M}_{k}$.
As shown in Appendix E.1, even if (1) is satisfied only at certain root states, one does not need to explore depths $N+1$ and beyond at these root states. Furthermore, Lemma E.1 establishes that the suboptimality bound $\epsilon_{N}$ decreases monotonically in NN.
@@ 5 Numerical Case Studies
In this section, we present numerical experiments that validate and complement the analytical results from prior sections. We also benchmark the proposed approaches against the state of the art.
@@
2-State 2-Action example: We consider a two-state, two-action baseline MDP, shown in Figure 2, and evaluate it under sensing costs (a) $k=0.01$ and (b) $k=0.25$.
@@
$\bullet$ Applying Theorem 4.1 yields a sensing cost threshold of $0.05$, making always sensing optimal in case(a).(a). Thus, the optimal policy for state 0 is $(R,sense)$, and for $1$ is $(B,sense)$.
@@
$\bullet$ For $k=0.25$, the optimal policy for both root states remains unchanged after $N\geq 2$. Indeed, the condition of Lemma 4.3 is satisfied at $N=2$. However, the condition of Theorem 4.4 is satisfied for $N=4$ which suggests that the optimal policy should remain unchanged for $N\geq 4;$ see Figure 3.
(The stronger lower bound on $V^{*}_{\mathcal{M}_{k}}$ used in Figure 3 is detailed in Appendix E.1.)
010.70.7, 110.30.3, 00.90.9, 00.20.2, 110.70.7, 00.30.3, 110.80.8, 110.10.1, 0
@@
Figure 2: A two-state two-action baseline MDP with actions $\{Red,Blue\}$ and $\alpha=0.5$.
Figure 3: Applying Thm 4.4 on MDP in Fig 2; $k=0.25$
@@
Experimental Setup. We benchmark the proposed SPI algorithm against the ATM heuristic (Krale et al., 2023), alongside widely used general-purpose POMDP planning algorithms, including SARSOP (Kurniawati et al., 2008), Fast Informed Bound (FIB) (Hauskrecht, 2000), using the POMDPs.jl framework (Egorov et al., 2017). Experiments are conducted across the ICU-Sepsis benchmark (Choudhary et al., 2024) and gridworld-based environments from Gymnasium (Towers et al., 2024), including Taxi and Frozen Lake. Algorithms such as Incremental Pruning (Cassandra et al., 1997) and PBVI (Pineau et al., 2003) exhibited prohibitive runtimes, failing to produce effective policies even for small state spaces, while other methods like QMDP (Littman et al., 1995) were consistently outperformed by our approach. Even popular online planning algorithms like PO-UCT (Silver and Veness, 2010) required significantly higher planning times to generate competitive policies.
These POMDP planners were initialized with a belief distribution corresponding to the starting state distribution, rather than a uniform distribution over all states. The experimental design, along with an analysis of POMCP’s performance, is provided in Appendix B.1. For SARSOP, the reward and sensing costs were scaled by 103 during policy computation to match performance on Frozen Lake.
ICU-Sepsis ($|\mathcal{S}|=716\ \&\ |A|=25$) : Choudhary et al. (2024) is a tabular MDP modeling personalized care for sepsis patients in the ICU, derived from electronic health record (EHR) data. The state space $\mathcal{S}$ models a patient’s health as a vector of clustered features (demographics, vital signs, body fluid levels) discretized into 716 states. The action space consists of discrete combinations of medication types and dosages. The agent receives 0 reward for non-terminal steps and +1 for survival ($\alpha=0.99$).
Frozen Lake ($|\mathcal{S}|=16/64\ \&\ |A|=4$) : The agent navigates across a frozen lake and receives a reward of +1+1 upon reaching the goal state ($\alpha=0.9$). In the slippery case, actions succeed with probability $\frac{1}{3}$, otherwise moving perpendicularly with equal probability $\frac{1}{3}$. We evaluate default 4x4, 8x8, and a custom, challenging 4x4 Frozen Lake grid (see Table 2). For Frozen Lake, we report expected reward over the initial state S and maximum computation times across sensing costs.
@@
Stochastic Taxi ($|\mathcal{S}|=500\ \&\ |A|=6$) : The agent navigates a 5x5 grid to pick up and drop off passengers, receiving rewards of +20 for delivery, -10 for illegal “pickup” and “drop-off” actions, and -1 per step unless another reward is triggered ($\alpha=0.95$(Dietterich, 1999). Each of the four navigation actions moves the agent in the intended direction with a probability of 0.8, and in one of the two perpendicular directions with an equal probability of 0.1. The initial states are sampled from 300 valid states.
Environment Sensing Costs
0.005 0.01 0.05 0.1
ICU-Sepsis
    $\pi_{\text{ATM}}$ Value 0.740 0.715 0.714 0.714
       Time (s) 285 782 833 916
SARSOP (3000s)
0.741 0.741 0.741 0.740
SPI (2 iter) Value 0.765 0.747 0.742 0.745
SPI (2 iter) Time (s) 23 1252 2205 2757
0.1 0.5 1 5
Stochastic Taxi
    $\pi_{\text{ATM}}$ Value 0.908 -0.359 -2.761 -19.664
       Time (s) 2 1.7 1.8 38.4
SARSOP (100s)
-22 -1.2 -2.705 -20
    SPI    Value 0.911 -0.306 -1.598 -9.778
    SPI    Time (s) 25.2 19.7 17.8 117.8
Scenario Value (Rewards) for Sensing Costs Time (s)
Frozen Lake 0.001 0.005 0.01 0.05
4x4 (Default)
   $\pi_{\text{ATM}}$ 62.42 36.52 6.72 16.57 0.04
   $V^{*}_{\mathcal{M}_{k,3}}$ 62.42 36.53 20.47 -28.75 11.5
   SARSOP 62.42 36.53 20.79 23.08 1.7
   FIB (1000 iter) 62.42 31.43 23.08 23.08 8.3
   SPI 62.42 36.53 20.99 23.08 0.4
4x4 (Hard)
   $\pi_{\text{ATM}}$ 8.41 0 0 0 0.04
@@
   SARSOP 8.95 3.66 1.34 1.44 1.6
   FIB (1000 iter) 4.44 0 0 0 9.5
   SPI 8.95 3.69 1.47 1.35 0.4
8x8
@@
Table 1: Performance comparison across different scenarios with varying sensing costs; rewards scaled by $10^{3}$ for Frozen Lake. (Unless the no. of $\mathsf{PolicyUpdate}$ iter in SPI are specified, $\delta=10^{-6}$.)
Benchmarking Results. Table 1 shows that SPI consistently outperforms $\pi_{\text{ATM}}$ and FIB, and achieves the best performance on ICU-Sepsis and Taxi across all sensing costs (FIB could not scale to these domains). Results from the custom hard setup further indicate that $\pi_{\text{ATM}}$ and FIB may fail to find effective policies even in small state spaces when the planning problem is challenging, underscoring SPI’s robustness.
We emphasize that, although SARSOP exhibited performance comparable to SPI for Frozen-Lake, achieving this required scaling the reward and sensing values. Without such scaling, SARSOP performed poorly, even relative to $\pi_{\text{ATM}}$ and FIB. Furthermore, in Stochastic Taxi, SARSOP consistently underperformed, and scaling the values did not improve its performance. Finally, we note that SARSOP’s performance is highly sensitive to several hyperparameters that are difficult to interpret and tune. In contrast, SPI relies on only two hyperparameters, both of which are intuitive and easily adjustable. Moreover, SARSOP requires the starting state distribution to compute its policy and can degrade significantly under mismatch, whereas SPI operates without requiring such information.
Finally, our analysis reveals significant limitations in existing RL algorithms for this setting, notably the Observe-before-planning (Nam et al., 2021) and Dyna-ATMQ (Krale et al., 2023). Observe-before-Planning leverages the ACNO-MDP structure during exploration but resorts to the generic POMDP solver POMCP (PO-UCT) for planning, failing to utilize the specific ACNO-MDP structure during exploitation. Furthermore, PO-UCT performs significantly worse than SPI, especially in scenarios with sparse rewards. Similarly, Dyna-ATMQ employs a less effective planner $\pi_{\text{ATM}}$, which struggles even in small, challenging state spaces. We attribute these inefficiencies to the absence of a strong theoretical foundation for ACNO-MDPs, a gap our paper addresses. By reformulating this setting using an “expanded state-space setup,” we provide novel insights that enable the design of effective planning algorithms like SPI.
6 Discussion and Conclusion